# RealmDreamer: Text-Driven 3D Scene Generation with Inpainting and Depth Diffusion

### v1

@@ 4.3 Depth diffusion for text-to-3D
As mentioned in the previous section, during training, we can draw samples $\hat{x}$ using the inpainting diffusion model $\epsilon_{\text{inpaint}}$ applied to a noisy rendering of the current scene. Our insight is to use these clean samples as the conditioning for the depth diffusion model (shown in Fig. 4). Starting from pure noise $d_{1}\sim\mathcal{N}(0,I)$, we predict the normalized depth using DDIM sampling [59]. We then compute the (negated) Pearson Correlation between the rendered depth and sampled depth:
@@
where $d$ is the rendered depth from the 3DGS model and $n$ is the number of pixels.
4.4 Optimization and Finetuning
@@
After training with this loss, we have a 3D scene that roughly corresponds to the text prompt, but which may lack cohesiveness between the reference image $I_{ref}$ and the inpainted regions. To remedy this, we incorporate an additional lightweight finetuning phase. In this phase, we utilize a vanilla text-to-image diffusion model $\epsilon_{\text{text}}$ personalized for the input image [52, 48, 15, 37]. We compute $\hat{x}$ using the same procedure as in Sec. 4.2, except with $\epsilon_{\text{text}}$. The loss $L_{\text{text}}$ is the same as Eq. 5, except with the $\hat{z}$ and $\hat{x}$ sampled with this finetuned diffusion model $\epsilon_{\text{text}}$.
To encourage sharp details in our model, we use a lower noise strength than the inpainting stage and uniformly sample the timestep $t$ within this limited range. We also propose a novel sharpening procedure, which improves the sharpness of our final 3D model. Instead of using $\hat{x}$ to compute the image-space diffusion loss introduced earlier, we use $\mathcal{S}(\hat{x})$, where $\mathcal{S}$ is a sharpening filter applied on samples from the diffusion model. Finally, to encourage high opacity points in our 3DGS model, we also add an opacity loss $L_{opacity}$ per point that encourages the opacity of points to reach either 0 or 1, inspired by the transmittance regularizer used in Plenoxels [16].
The combined loss for the fine-tuning stage can be written as:
$L_{\text{finetune}}=L_{\text{text}}+\lambda_{opacity}L_{\text{opacity}},$ (8)
where $\lambda_{opacity}$ is a hyperparameter controlling the effect of the opacity loss.
Figure 6: We show the main qualitative results from our proposed method. In the left column, we show the input prompt. In the next two columns, we show the renderings from our 3D model from different viewpoints. In the fourth column, we show the level of agreement between rendering and geometry by a split view of the rendering and depth. Finally, in the last column, we show the depth map.
@@ 4.5 Implementation Details
Point Cloud Initialization. We implement the point cloud initialization stage (Sec. 4.1) in Pytorch3D [49], with Stable Diffusion [51] as our inpainting model. To lift the generated samples to 3D, we use a high quality monocular depth estimation model, Marigold [29], as a relative depth estimator. Since it lacks a defined scale, we align these predictions with the absolute metric depth predicted with DepthAnything [69].
Inpainting and Finetuning Stage. Our inpainting (Sec. 4.2) and finetuning stages (Sec. 4.4) are implemented in NeRFStudio [62] using the official implementation of Gaussian Splatting [30]. We use Stable Diffusion 2.0 as $\epsilon_{\text{text}}$ and its inpainting variant as $\epsilon_{\text{inpaint}}$, building on threestudio [18] to define our diffusion-guided losses. Further, we use Marigold [29] as our depth diffusion model. During the inpainting stage, we set the guidance weight for image and text conditioning of $\epsilon_{\text{inpaint}}$ as 1.8 and 7.5 respectively. We find that a high image guidance weight produces samples with greater overall cohesion. We also use a guidance weight of 7.5 for the text-to-image diffusion model $\epsilon_{\text{text}}$ during the finetuning stage. During the inpainting stage, we sample the timestep $t$ from $\mathcal{U}(0.1,0.95)$ and during the finetuning stage from $\mathcal{U}(0.1,0.3)$. We train the model for 15000 iterations during the inpainting stage, which takes 8 hours to run on a 24GB Nvidia A10 GPU. The finetuning stage is run for 3000 iterations and takes 2.5 hours to finish on the same GPU.
Figure 7: We show the main qualitative comparison to the state-of-the-art baselines ProlificDreamer [66], DreamFusion [46], and Text2Room [23] on the two prompts specified.
Figure 8: We show the qualitative results of our model and its ablations. Arrows indicate failures in the ablated models. Please see  Sec. 5.5 for a detailed discussion of the ablated components and their respective importance.
@@ 5 Results
We evaluate our technique on a custom dataset of prompts, reference images, and associated camera poses $P_{i}$, selected to showcase parallax and disocclusion in the 3D scene. We built this dataset by creating a set of 11 prompts, and having a human expert manually choose camera poses using a web-viewer [62], by displaying a scene prototype obtained as in Sec. 4.1. No such dataset already exists for this problem, as existing text-3D techniques [46, 66] operate with simple camera priors, such as sampling on a sphere, with cameras looking at a central object, or outward in the case of scenes [66]. This can limit the scale of the obtained scene significantly.
@@ 5.1 Qualitative Results
Our method’s results are shown on 5 representative examples in Fig. 6, demonstrating effective 3D scene sampling across various settings (indoor, outdoor) and image styles (realistic, fantasy, illustration). We highlight the significant difference between views and the consistency of rendering and geometry in split views, underscoring our method’s use of inpainting and depth diffusion priors.
@@ 5.2 Comparisons
We compare our technique with the state-of-the-art baselines for text-to-3D synthesis in  Fig. 7. Namely, we compare against DreamFusion [46], ProlificDreamer [66], and Text2Room [23]. We choose these baselines as they represent the state-of-the-art performance for 3D object and scene synthesis. Both ProlificDreamer and DreamFusion learn oversaturated scenes with incorrect geometry and scene structure. On the other hand, Text2Room fails to reconstruct non-room scenes, resulting in degenerate renderings which essentially forget the input prompt during its iterative mesh creation process. To ensure a fair comparison, we also incorporate prompts from the baselines Text2Room [23], ProlificDreamer [66], and Dreamfusion [46].
@@ 5.3 User Study
In line with prior work [66], we conduct a user study to compare the results from our technique with the established baselines. We conducted the study on Amazon Mechanical Turk and recruited 20 participants with a ‘master’ qualification to compare with each baseline, while following several guidelines outlined in [5]. All participants were provided a brief definition about “3D scenes” and asked to make preferential decisions based on three criteria: 1). the quality of the videos 2). how 3D the scene looks and 3). alignment to the associated prompt.
For comparison with ProlificDreamer [66] and DreamFusion [46], we showed participants side-by-side videos comparing our method to the baseline. For fairness, we use the same camera trajectory in all videos. The order of the videos was also randomized to prevent any biases due to the order of presentation. The user’s preference was logged along with a brief explanation.
When comparing with Text2Room [23], instead of a video, we showed users side-by-side sets of three multiview images for each prompt, due to the degeneracy of the output mesh far from the starting camera pose. The users preferred triplet was logged along with their brief explanation. The images we showed looked slightly left and right of the reference pose $P_{\text{ref}}$.
Table 1: Results of user study. This table shows the percentage of comparisons where our technique was preferred over the mentioned baseline. Our approach is significantly preferred over existing baselines.
Ours vs. ProlificDreamer[66] Ours vs. DreamFusion[46] Ours vs. Text2Room[23]
93.63% 92.27% 87.72%
5.3.1 Common Themes
ProlificDreamer [66] can produce cloudy results. Several participants described the NeRF renders as containing “moving clouds”, a “hazy atmosphere”, and a “blotch of colours”. This can likely be attributed to the presence of floaters in the model, which is evident in the noisy depth maps shown in Sec. A.2. In contrast, participants described our method as “clean and crisp when it comes to the colors and sharpness of the pixels” and looking realistic, without the presence of over-saturated colors.
@@ 5.4 Quantitative Metrics
We also provide a quantitative comparison between our method and baselines using CLIP [47]. Specifically, for each method, we compute the CLIP alignment between the rendered image at $P_{\text{ref}}$ and the corresponding prompt for all scenes. For Text2Room, this pose is the input pose from which the scene is generated. The results can be seen in Tab. 2, where our method shows better alignment to the text prompt versus the baselines. Note that this is due to our method’s ability to inherit the high-fidelity generations of 2D diffusion models.
Ours
Text2Room
DreamFusion
ProlificDreamer
33.14
25.10
30.28
  30.69
Table 2: CLIP alignment for renderings of scenes for our method and the baselines. Scores are scaled by 100.
@@ 5.5 Ablations
We verify the proposed contributions of our method by ablating the key components in  Fig. 8 with the specified prompt. In the first row, we show our base method. In the second row, we then show the importance of the inpainting diffusion model from Sec. 4.2—removing this component results in high-variance gradients which cause the 3DGS representation to diverge into a meaningless cloud. In the third row, we remove the depth prior and $L_{\text{depth}}$; this results in a 3DGS representation with incorrect geometry which doesn’t reflect the renderings. Next, in the fourth row, we initialize the point cloud with only our reference image $I_{ref}$, without sampling from neighbouring poses $P_{\text{aux}}$, which makes the regions outside of the original camera frustum underdetermined and low quality as evidenced by the images. Finally, in the last row, we show our result without applying a sharpening filter on the sampled images which results in blurrier renders with less detail.
6 Application: Single image to 3D
@@
Our technique extends to creating 3D scenes from a single image, as shown in Fig. 9, by using a user’s image as $I_{ref}$ and a text-prompt $T_{ref}$ obtained using an image-captioning model like GPT-4. Our pipeline remains the same and can effectively fill in occluded areas and generate realistic geometry for unseen regions.
@@ 7 Conclusion
We have proposed RealmDreamer, a method for generation of forward-facing 3DGS scenes leveraging inpainting and depth diffusion priors. Our method provides higher-quality results than existing baselines along with higher-fidelity geometry on more challenging camera trajectories. Still, some limitations remain; our method takes several hours, and produces blurry results for complex scenes with high occlusions. Future work may explore more efficient diffusion models for faster training, and more complex conditioning schemes for full 360-degree generations.
@@
8 Acknowledgements
Appendix A Appendix
In this supplementary, we provide additional details concerning RealmDreamer, such as additional implementation details (Sec. A.1) and additional qualitative results (Sec. A.2).
A.1 Additional Implementation Details

### v2

@@ 4.3 Depth Diffusion for Geometry Distillation
Our insight is that these samples $\hat{x}$ act as suitable, in-domain, conditioning for the depth diffusion model throughout optimization, while renders $x$ can be incoherent before convergence. Further, this ensures that predictions from $\epsilon_{\text{depth}}$ are aligned with $\epsilon_{\text{inpaint}}$ despite not using a RGBD prior. Starting from pure noise $d_{1}\sim\mathcal{N}(0,I)$, we predict the normalized depth using DDIM sampling [59]. We then compute the (negated) Pearson Correlation between the rendered depth and sampled depth:
@@
where $d$ is the rendered depth and $n$ is the number of pixels.
4.4 Optimization and Refinement
@@
After training with this loss, we have a 3D scene that roughly corresponds to the text prompt, but which may lack cohesiveness between the reference image $I_{ref}$ and the inpainted regions (see Fig. 4). To remedy this, we incorporate an additional lightweight refinement phase. In this phase, we utilize a vanilla text-to-image diffusion model $\epsilon_{\text{text}}$ personalized for the input image with Dreambooth [51, 47, 17, 39]. We compute $\hat{x}$ using the same procedure as in Sec. 4.2, except with $\epsilon_{\text{text}}$. The loss $L_{\text{text}}$ is the same as Eq. 5, except with the $\hat{z}$ and $\hat{x}$ sampled with this finetuned diffusion model $\epsilon_{\text{text}}$. Note that the noise added to the renderings at this stage is smaller to combat the higher variance samples from the lack of image conditioning.
We also propose a novel sharpening procedure: instead of using $\hat{x}$ to compute the image-space diffusion loss introduced earlier, we use $\mathcal{S}(\hat{x})$, where $\mathcal{S}$ is a sharpening filter applied on samples from the diffusion model. Finally, to encourage high opacity points in our 3DGS model, we incorporate an opacity loss $L_{\text{opacity}}$ per point that encourages a point’s opacity to reach either 0 or 1, inspired by the transmittance regularizer used in Plenoxels [19]. The combined loss for the fine-tuning stage is:
$L_{\text{refine}}=L_{\text{text}}+\lambda_{\text{opacity}}L_{\text{opacity}},$ (8)
where $\lambda_{\text{opacity}}$ controls the effect of the opacity loss.
@@ 4.5 Implementation Details
Point Cloud Initialization. We implement this stage (Sec. 4.1) in Pytorch3D [48], with Stable Diffusion [50] for outpainting. To lift the generated images to 3D, we use Marigold [30], a monocular depth estimation model. Since it predicts relative depth, we align its predictions with the metric depth predicted by DepthAnything [65].
Inpainting and Refinement Stage. Our inpainting (Sec. 4.2) and refinement stages (Sec. 4.4) are implemented in NeRFStudio [62] using the official implementation of Gaussian Splatting [31]. We use Stable Diffusion 2.0 as $\epsilon_{\text{text}}$ and its inpainting variant as $\epsilon_{\text{inpaint}}$, building on threestudio [22] to define our diffusion-guided losses. Further, we use Marigold [30] as our depth diffusion model. During the inpainting stage, we set the guidance weight for image and text conditioning of $\epsilon_{\text{inpaint}}$ as 1.8 and 7.5 respectively, and sample the timestep $t$ from $\mathcal{U}(0.1,0.95)$. We find that a high image guidance weight produces samples with greater overall cohesion. We also use a guidance weight of 7.5 for the text-to-image diffusion model $\epsilon_{\text{text}}$ during the refinement stage, sampling noise from $\mathcal{U}(0.1,0.3)$.
Timing. The first stage, currently unoptimized, takes 2.5 hours. The inpainting stage, trained for 15,000 iterations, runs for 8 hours on a 24GB Nvidia A10 GPU. The refinement stage, at 3,000 iterations, completes in 2.5 hours on the same GPU.
@@ 5 Results
We evaluate our technique on a custom dataset of 20 prompts, and associated camera poses $P_{i}$, selected to showcase parallax and disocclusion. We built this dataset by creating a set of 20 prompts, and having a human expert manually choose camera poses using a web-viewer [62], by displaying a scene prototype obtained as in Sec. 4.1. No such dataset already exists for this problem, as existing text-to-3D techniques [44, 64] typically operate with spherical camera priors. Please refer to the supplemental video results to see the generated scenes.
@@ 5.1 Qualitative Results
We show some qualitative results in Fig. 5 with additional results in the supplementary, demonstrating effective 3D scene synthesis across various settings (indoor, outdoor) and image styles (realistic, fantasy, illustration). We would like to highlight the rendering quality and the consistency of rendering and geometry, underscoring our method’s use of inpainting and depth priors.
Figure 7: Ablation Results. We show the qualitative results of our model and its ablations. Arrows indicate failures in the ablated models. Please see  Sec. 5.5 for a detailed discussion of the ablated components and their respective importance.
@@ 5.2 Comparisons
We compare our technique with state-of-the-art for text-to-3D that use either distillation or iterative approaches: DreamFusion [44], ProlificDreamer [64], Text2Room [26], and concurrent work LucidDreamer [14] (Fig. 6). Both ProlificDreamer and DreamFusion generate oversaturated scenes with incorrect geometry and scene structure. On the other hand, Text2Room fails to construct non-room scenes, as it deviates from the input prompt during generation. Similarly, LucidDreamer’s [14] scenes lack cohesion, with noisy results in occluded regions. Note that LucidDreamer and Text2Room take an image as input; we gave these baselines the same input image as to ours and updated their depth models.
@@ 5.3 User Study
To validate the quality of our generated 3D scenes, we conduct a user study (Tab. 1), similar to prior work [64, 35, 12]. We conducted a study on Amazon Mechanical Turk and recruited 20 participants with a ‘master’ qualification to compare each baseline, while following several guidelines outlined in [7]. The details can be found in the supplementary. Participants overwhelmingly prefer results from our technique over baselines.
Table 1: Results of user study. We show the percentage of comparisons where our technique was preferred over baselines: PD [64], DF [44], T2R [26], and LD [14].
Ours vs. PD Ours vs. DF Ours vs. T2R Ours vs. LD
95.5% 94.5% 88% 88.5%
Table 2: CLIP alignment scores and additional metrics for scene renderings of our method and the baselines. CLIP scores are scaled by 100. Higher is better for all metrics.
Method CLIP Depth Pearson IS
Ours 31.69 0.89 6.99
Text2Room [26] 28.11 0.77 5.10
DreamFusion [44] 29.48 0.09 6.80
ProlificDreamer [64] 29.39 0.16 6.89
LucidDreamer [14] 29.97 0.80 5.73
@@ 5.4 Quantitative Metrics
We also provide a quantitative comparisons with all baselines based on alignment to the text prompt using CLIP [46], Inception Score [52] on renderings, and the quality of geometry with the pearson correlation between rendered depth and the predicted depth by DepthAnythingV2 [66]. We note that due to the lack of ground truth data, standard reconstruction metrics such as PSNR or LPIPS [69] do not apply. We compute these scores for renderings from the same trajectory and the corresponding prompt for all scenes. As Text2Room’s results degrade significantly away from the initial pose, we compare with a render from the initial pose for CLIP. As shown in Tab. 2, our method shows significantly better performance across all metrics.
@@ 5.5 Ablations
We verify the proposed contributions of our method by ablating the key components in  Fig. 7 with the specified prompt (Tab. 3). In the first row, we show our method. In the second row, we show the importance of the low variance samples from the inpainting diffusion model (Sec. 4.2). Distillation with a vanilla text-to-image model as in the final stage, results in high-variance samples causing the 3DGS representation to diverge. In the third row, we remove $L_{\text{depth}}$; this results in incorrect geometry and incoherent renderings. Note in particular the discrepancy in the background when viewing from left versus right. In the fourth row, we initialize our method using only the reference image $I_{ref}$ without outpainting at the neighbouring poses $P_{\text{aux}}$. This results in poor results in the corresponding regions, as they lack a good initialization. Finally, in the last row, we show our result without using the $\mu$ initialization from Eq. 4, which results in divergence.
Table 3: Ablation Study Results showing the impact of different components on Depth Pearson correlation and CLIP score. CLIP scores are scaled by 100. Higher is better for both metrics.
Ablation Depth CLIP
No Depth Loss 0.86 31.55
No Initialization 0.42 20.31
No Inpainting 0.50 21.14
No Outpainting 0.79 31.00
Ours 0.90 33.10
5.6 Application: Single image to 3D
@@
Our technique extends to creating 3D scenes from a single image, as shown in Fig. 8, by using a user’s image as $I_{ref}$ and a text-prompt $T_{ref}$ obtained using an image-captioning model. Our pipeline can effectively fill in occluded areas and generate realistic geometry for unseen regions.
@@ 6 Conclusion
We have proposed RealmDreamer, a method for generation of forward-facing 3DGS scenes leveraging inpainting and depth diffusion. Our key insight was to leverage the lower variance of image conditioned (inpainting) diffusion models for synthesis of 3D scenes, providing much higher quality results than existing baselines as measured by a comprehensive user study. Still, limitations remain; our method takes several hours, and produces blurry results for complex scenes with significant disocclusion. Future work may explore efficient diffusion models for faster training, and conditioning for 360-degree generations.
7 Acknowledgements
Supplemental Material:
RealmDreamer: Text-Driven 3D Scene Generation
with Inpainting and Depth Diffusion
Figure 9: Creating the Inpainting Mask. 3D inpainting requires filling in a 3D volume, which is not always equivalent to missing regions in 2D point cloud renders. By computing an occlusion volume, we avoid situations where the floor is visible through the table (middle) but instead could be occluded. The right depth map accounts for the ambiguity of the volume in the table.
@@ 7 Acknowledgements
Appendix A Why is the occlusion volume important?
Our key contribution is the inpainting distillation loss that provides lower variance and high-quality supervision for text-to-3D, compared to regular text-to-image model based distillation, as shown in the ablations. Given that we use 2D inpainting models for 3D inpainting via this distillation, we must ask: how to compute the 3D region that needs to be inpainted?
We proposed a simple technique to do so by computing the occluded volume (described in Sec. D.2), which is the 3D region occluded by objects in the reference image $I_{\text{ref}}$. Note that regardless of what objects may be present in this occluded region, rendering from $P_{\text{ref}}$ would yield $I_{\text{ref}}$, as elements in the image would occlude the objects. The 2D inpainting masks we obtain then must hence, reflect this unknown 3D region, as that is the part of the scene left to complete.
Instead of computing the occlusion volume, another alternative is using the holes from point cloud renderings as the inpainting mask. Indeed, these holes also represent unknown 3D regions. However, the 3D region indicated by such 2D masks is a subset of the occluded 3D region and hence, incomplete. This is shown in Fig. 9, which shows the masked point cloud depth, where the masked region represents the inpainting mask. When using the holes in the point cloud as an inpainting mask, one can observe that the back side of the kitchen table is visible. In reality, a solid kitchen table would never expose this face. In contrast, using the occlusion volume, we can correctly determine the entire 3D region missing in $I_{\text{ref}}$, which can be visually verified by comparing images. Specifically, the latter takes into account self-occlusion, providing a more accurate estimate and allowing details to fill into the occluded region.
In practice, such self-occlusions do not appear for all prompts yet is important to maintain the correctness of the 3D inpainting formulation. If there are many renderings such as in Fig. 9, the quality of inpainted samples would be incorrect and lead to a noisier distillation process.
Appendix B Discussion on Baselines
We are among the first to tackle text-to-3D scenes and showcase a high-level of parallax. As a result, there are limited open-source 3D scene generation techniques to compare with. We choose to compare with techniques that either use distillation - a key part of our pipeline and iterative approaches - which shares similarity with the initialization technique we use.
B.1 Comparison with Dreamfusion and ProlificDreamer
By comparing with prior SOTA distillation techniques [44, 64], we demonstrate that these approaches are suboptimal for scene generation, which has not been demonstrated before. Arguably, distillation techniques should be able to build any 3D scene since the 2D priors used are general, assuming object-centric regularizers and prompts are absent. However, we empirically find that simple distillation from text-to-image models is insufficient for wide baselines. This comparison highlights the importance of our inpainting distillation, which conditions on a partial 3D scene, enabling high level of parallax not seen in prior work.
We also note that ProlificDreamer [64] first showcased scene-level results using distillation, indicating that distillation is not purely for object-generation. Still, it presented a limited set of scenes and did not show high parallax as it focused on rotating a camera through a scene. Our baseline comparison attempts to test the performance of this approach on wide camera trajectories and finds that it often produces hazy results.
B.2 Relation between SDS and our distillation
Our distillation is similar to the score distillation loss (SDS) used in Dreamfusion [44]. However, unlike prior work, we do not use just text conditioning, but also renders from the point cloud. As a result, the classifier-free guidance weight we is much lower - at 7.5, avoiding over-saturated results typically found with SDS. Further, unlike SDS which denoises noisy renders in one step, we take multiple steps with DDIM sampling. Further, we also use a loss on the denoised latent and decoded image, to produce high quality supervision.
B.3 Comparison with LucidDreamer and Text2Room
Our initialization step is a key part of the pipeline and reminiscent of prior and concurrent iterative techniques which incrementally grow a scene. Yet, such techniques have yet to showcase high quality over high-parallax camera trajectories, which we demonstrate. We find that incremental generation of 3D scenes can lead to noise accumulation, due to errors in monocular depth and alignment of geometry. Hence, unlike concurrent work LucidDreamer [14], we do not limit our scene generation to our initialization step. We rely on our inpainting distillation loss to produce highly cohesive 3D scenes, by distilling across multiple views, rather than building a scene incrementally. We also note that while LucidDreamer does use 3DGS optimization, it is a more conventional reconstruction-based optimization, with no inpainting or geometry priors incorporated. As a result, its optimization stage is very different from our inpainting distillation, which provides rich priors for appearance and geometry at every iteration.
Further, we also avoid many limitations of prior work such as Text2Room, which often produces scenes with low-prompt alignment, especially those in outdoor scenes. We attribute this to the technique deleting regions of the mesh before inpainting, as the original technique prescribes. When such deletions accumulate over time, they are prone to erasing parts of the original scene defined in $I_{\text{ref}}$ entirely. In contrast to this, our technique maintains a simple initialization strategy and relies on a high-quality inpainting distillation process to fill-in mission regions, without sacrificing the quality of the initialized regions.
B.4 Implementation
Text2Room [26] and LucidDreamer [14] We use the official implementation of Text2Room and LucidDreamer on Github. To ensure a fair comparison, we estimate depth using Marigold [30] and DepthAnything [65] as in our technique, replacing the original IronDepth [3] and ZoeDepth [5] respectively. The rest of the pipeline is kept the same.
Appendix C Additional ablations and discussion
C.1 Use of DDIM Inversion.
During the inpainting and refinement stage (Sec 4.2, 4.3 in the original paper), we find it helpful to obtain the noisy latent $z_{t}$ using DDIM inversion [59], where $z=\mathcal{E}(x)$, $x$ is the rendered image, and $t$ is a timestep corresponding to the amount of noise added. This is similar to prior work on 2D/3D editing and synthesis using pre-trained diffusion models [23, 33]. We demonstrate the importance of doing so in Fig. 10, where DDIM inversion can significantly improve the detail in the optimized model. During the inpainting stage, we use 25 steps to sample an image from pure noise, and during refinement, we use 100 steps.
C.2 Use of sharpening filter.
In Fig. 10, we also see that applying a sharpening filter to the sampled images results in slightly more detail. We attribute this to the blurry nature of some samples of the diffusion model.
Figure 10: We ablate the importance of DDIM Inversion and applying a sharpening filter. As in  [33], we find that DDIM inversion allows more details to be synthesized by our method. Additionally, we find that detail slighly increases when applying a sharpening filter to the sampled images.
Appendix D Additional Implementation Details